OpenStudy (anonymous):

The midpoint of CD is E(-1, 0). One endpoint is C(5, 2). What are the coordinates of the other endpoint? Show all work.

OpenStudy (anonymous):

Assume the other endpoint is D with coordinates (x,y) Because E is midpoint of CD, we have the equation: \[\frac{x+5}{2} = -1\] \[\frac{y+2}{2}=0\] Solve for x, y, we obtain: x = -7, y = -2. So coordinates of D is (-7,-2)

OpenStudy (anonymous):